# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v1

@@ 6 Conclusion
This paper presented a data-driven contextual stochastic optimization framework, which was applied to an industrial omnichannel order fulfillment problem within the context of multi-courier logistics. Leveraging contextual optimization effectively addresses the inherent complexities of modern retail logistics, providing innovative solutions tailored to the dynamic e-commerce landscape. In particular, the proposed framework is capable of handling the availabilities of observational data only.
To achieve these results, the paper uses probabilistic multi-class classification and quantile regression models to estimate the conditional distribution of delivery time deviations. These machine learning models are then seamlessly integrated with SAA-based risk-neutral optimization models and robust optimization models under the proposed CSO framework. The paper evaluated the proposed methodology on a real dataset from one of the largest e-commerce company. The results, which significantly improve the baseline solution, showcase the ability of the CSO framework to improve service quality, customer satisfaction, and operational efficiency. They offer valuable insights and tools for practitioners aiming to optimize their order fulfillment strategies in the evolving e-commerce environment.
There are several potential directions for future work. The contextual optimization models could be extended to incorporate long-term inventory planning through multi-period order fulfillment optimization. This would shift the approach from a rather myopic perspective to a more comprehensive strategy that accounts for future inventory needs. The distributional forecasts of deviations could be improved by increasing the dataset to include factors such as seasonal trends, weather conditions, and carrier tracking details. Furthermore, integrating downstream optimization costs into the training of distributional forecasting models can be an interesting extension (Qi et al. 2021, Elmachtoub and Grigas 2022). However, it is crucial to address the observational nature of the dataset and the varying dimensions of different optimization problem instances. Another promising direction is to extend the current contextual optimization framework to an end-to-end optimization approach (Kotary et al. 2021) while adhering to all operational constraints .
This paper is relevant to the broader literature on uncertainty quantification, where conformal predictions offer a promising direction. Conformal predictions provide coverage guarantees for prediction intervals and could be particularly effective when combined with quantile regression (Romano et al. 2019). Moreover, adopting conformally calibrated predictive distributions (Manokhin 2017) and improving probability calibration methods could enhance model performance. The paper hints at this potential by adopting probability calibration in multi-class classification. The Venn-Abers calibrator was experimented, but it was too computationally intensive to be fitted for the dataset used in the paper.
[…]
This research was partly supported by the NSF AI Institute for Advances in Optimization (Award 2112533).

### v2

@@ 7 Conclusion
This paper introduces a generic contextual stochastic optimization (CSO) framework for data‑driven decision making when only observational data—hence missing counterfactual costs—are available. The framework is deliberately modular. First, any algorithm capable of returning a full predictive distribution can serve as the contextual distribution oracle. In the empirical study, calibrated multi‑class classifiers and tree‑based quantile‑regression forests were selected to capture the discrete, ordered nature of delivery‑time deviations. Second, the oracle feeds directly into two tractable optimization engines: (i) a Contextual Sample‑Average Approximation (C‑SAA) for risk‑neutral objectives and (ii) a Contextual Robust Optimization (C‑RO) formulation for risk‑averse objectives. Because the learning and optimization layers communicate only through the oracle, the template can be readily translated to other operations‐management settings.
When applied to an omnichannel, multi‑courier order‑fulfillment problem, the CSO framework reduced the combined cost–service objective by up 18% on average and up to 22% in the simulated worst case on an industry‑scale data set, implying annual savings from several to tens of millions of dollars. This framework allows practitioners to tune robustness—raising it during promotional peaks and relaxing it in steadier periods—while still exploiting item‑consolidation discounts and heterogeneous carrier performance in the dynamic e-commerce landscape.
While the paper delivers an end‑to‑end prototype with demonstrable impact, several avenues remain open:
  • Performance guarantees under partial feedback: Establish theoretical performance guarantees, such as worst-case regret and post-decision surprise, for CSO methods under the partial‑feedback setting. Linking causal‑inference techniques with SAA and RO theory can be a promising direction, thereby quantifying the value of additional data when counterfactual costs remain unobserved.
  • Hybrid distributional forecasters: Develop distributional models that fuse the calibration strength of classification with the sharp tails of quantile regression, yielding tighter and more informative uncertainty sets.
  • End‑to‑end predict‑then‑prescribe learning for observational data: Incorporate downstream optimization loss directly into training, extending recent “integrated predict‑then‑optimize” approaches (Qi et al. 2021, Elmachtoub and Grigas 2022) to settings where data are purely observational.
  • Multi‑period, inventory‑aware CSO: Extend the framework to jointly optimize sourcing, shipping, and stocking over rolling horizons, capturing inventory replenishment spillovers and elevate the model from a myopic to a long-term strategic planning tool.